OpenStudy (anonymous):

how do i find the arc of a circle of radius 81cm intercepted by the central angle of a=2radians???

OpenStudy (anonymous):

please help

jimthompson5910 (jim_thompson5910):

what's the circumference of this circle

OpenStudy (anonymous):

not given

jimthompson5910 (jim_thompson5910):

can you find the circumference of this circle

OpenStudy (anonymous):

2 pie r

jimthompson5910 (jim_thompson5910):

r is what

OpenStudy (anonymous):

81

jimthompson5910 (jim_thompson5910):

so it's 2*pi*81 = ???

OpenStudy (anonymous):

calculater says error?

jimthompson5910 (jim_thompson5910):

replace pi with 3.14

OpenStudy (anonymous):

ok so 508.68

OpenStudy (anonymous):

arc / circumfrence = angle / 2 pie?>

jimthompson5910 (jim_thompson5910):

now compute 2/(2pi) = 2/(2*3.13) = ???

jimthompson5910 (jim_thompson5910):

i mean 2/(2pi) = 2/(2*3.14) = ???

OpenStudy (anonymous):

???? those aret he same is that all it is

jimthompson5910 (jim_thompson5910):

what is 2/(2*3.14)

OpenStudy (anonymous):

0.318

jimthompson5910 (jim_thompson5910):

that's the number you multiply with the with circumference to get your answer

jimthompson5910 (jim_thompson5910):

0.318*(Circumference) = ???

OpenStudy (anonymous):

161.76

OpenStudy (anonymous):

so redius = pie * circumfrence>?

OpenStudy (anonymous):

or 2 pie?

OpenStudy (anonymous):

circum is 2pie r and r=2 pie ?

jimthompson5910 (jim_thompson5910):

not sure what you're saying

OpenStudy (anonymous):

2/(2pi) = 2/(2*3.14) = ???

OpenStudy (anonymous):

i got 0.318 from that.. how was that equation determined?

jimthompson5910 (jim_thompson5910):

(given angle)/(number of radians in a complete circle)

OpenStudy (anonymous):

2/ (2pie) is given angle and whats the 2 pie?

jimthompson5910 (jim_thompson5910):

2pi is the number of radians in a complete circle

OpenStudy (anonymous):

nvm gor ir thx!!